# Customizing Text-to-Image Diffusion with Camera Viewpoint Control

### v1

@@ 3.2 Customization with Camera Pose Control
where $M$ is the object mask and $\lambda_{i}$ are hyperparameters to control the rendering quality of intermediate images vs the final denoised image. We keep $\lambda_{i}$ fixed across all experiments. We assume access to the object’s mask in the image, which is used to calculate the above losses. The three losses corresponding to FeatureNeRF are averaged across all pose-conditioned transformer layers.
Figure 4: Qualitative comparison. Given a particular target pose, we show the qualitative comparison of our method with (1) Image editing methods SDEdit, InstructPix2Pix, and LEDITS++ which edit a NeRF rendered image from the input pose, (2) Vica-NeRF, a 3D editing method that trains a NeRF model for each input prompt, and (3) LoRA + Camera pose, our proposed baseline where we concatenate camera pose information to text embeddings during LoRA fine-tuning. Our method performs on par or better in keeping the target identity and poses while incorporating the new text prompt—e.g., putting a picnic table next to the SUV car ($1^{\text{st}}$ column)—and following multiple text conditions—e.g., turning the chair red and placing it in a white room ($3^{\text{rd}}$ column). V token is used only in ours and the LoRA + Camera pose method. Ground truth rendering from the given pose is shown as an inset in the first three rows. We show more sample comparisons in Figure 15 of Appendix.
Figure 5: Qualitative samples with varying pose. Our method’s results with different text prompts and target poses as conditions. Our method learns the identity of custom objects while allowing the user to control the camera pose and text prompt for generating the object in new contexts, e.g., changing the background scene or object color and shape. In each row, the images were generated with the same seed while changing the camera pose around the object in a turntable manner. Figure 16 in the Appendix shows more such samples. Note that each image in a row is independently generated. We do not aim to generate multi-view consistent scenes.
@@
Training details.   During training, we sample the $N$ views equidistant from each other and use the first as the target viewpoint and the others as references. We modify $12$ transformer layers with pose conditioning out of $70$ transformer layers in Stable Diffusion-XL. For rendering, we sample $24$ points along the ray. The new concept is described as “V category”, with V as a trainable token embedding [45, 23]. Furthermore, to reduce overfitting [74], we use generated images of the same category, such as random car images with ChatGPT-generated captions [12]. These images are randomly sampled $25\%$ of the time during training. We also drop the text prompt with $10\%$ probability to be able to use classifier-free guidance. We provide more implementation details in Appendix C.
@@ 4 Experiments
Dataset.   For our experiments, we select concepts from the Common Objects in 3D (CO3Dv2) dataset [71], commonly used for novel view synthesis, and NAVI [33]. Specifically, we select four categories with three instances from the CO3Dv2 dataset- car, chair, teddy bear, and motorcycle—as each instance is uniquely identifiable for these categories. From the NAVI dataset, we select two unique toy-like concepts. We use the camera pose provided in the dataset for the multi-view images. A representative image of each concept is shown in Figure 14, Appendix A. For each instance, we sample $\sim 100$ images and use half for training and half for evaluation. The camera poses are normalized such that the mean of camera location is the origin, and the first camera is at unit norm [100].
@@
Table 1: Human preference evaluation. Our method is preferred over all baselines for text alignment, image alignment to target concept, and photorealism except LoRA + Camera pose, which overfits the training images, as also shown in Figure 4.
@@
Table 2: Camera pose accuracy in generated images by ours and the LoRA + Camera pose baseline method. We observe that the baseline usually overfits to training images and does not respect the target pose with new text prompts.
Baselines.   While no prior method targets our exact task, we use three types of related baselines – (1) 2D image editing methods, which aim to preserve details of the input image and thus keep the object in the same pose as the input image. This includes three recent and publicly available methods: LEDITS++ [8], InstructPix2Pix [9], and SDEdit [52] with Stable Diffusion -1.5 (and SDXL in Appendix A). As image editing methods by themselves do not support camera viewpoint manipulation, we first render a NeRF model [87] in the target pose and then edit the rendered image. (2) Customization-based method, LoRA+Camera pose, where we modify LoRA [76, 32] by concatenating the camera pose information to the text embeddings, following recent work Zero-1-to-3 [49]. (3) VICA-NeRF [20], a 3D editing method that trains a NeRF for each new text prompt. In Appendix C, we provide more details on implementation and hyperparameters for each baseline.
Evaluation metrics.   To create an evaluation set, we generate $16$ prompts per object category using ChatGPT [12]. We instruct ChatGPT to propose four types of prompts: scene change, color change, object composition, and shape change. We then manually inspect them to remove implausible or overly complicated text prompts [93]. Table 5 in Appendix B lists all the evaluation prompts.
For a quantitative comparison, we primarily use a pairwise human preference study. We compare our method against each baseline to measure image alignment to target concept, alignment to input text prompt, and photorealism of generated images. In total, we collect $\sim 1000$ responses per pairwise study using Amazon Mechanical Turk. We also show the performance of our method and baselines on other standard metrics like CLIP Score [67] and DINOv2 [59] image similarity [74] to measure the text- and image-alignment.
To measure whether the object in generated images corresponds to the input camera pose for our method and the LoRA + Camera pose baseline, we use a pretrained model, RayDiffusion [101], to predict the poses from generated images and calculate its error relative to the ground truth camera poses. More details about evaluation are provided in Appendix B.
@@ 4.1 Results
Figure 7: Extrapolating camera pose from training views. Our method can generalize to different camera poses, including viewpoints not within the training distribution. Top left: We vary the focal length from ×$\times 0.8$ to ×$\times 1.4$ of the original focal length. Top right: We vary the camera position towards the image plane along the $z$ axis. Bottom row: We vary the camera position along the horizontal and vertical axis.
@@
Generation quality and adherence.   First, we measure the quality of the generation – adherence to the text prompt, the identity preservation to the customized objects, and photorealism – irrespective of the camera pose. For the comparison, we generate $18$ images per prompt on $6$ target camera poses, totaling $288$ images per concept. Table 1 shows the pairwise human preference for our method vs. baselines. Our method is preferred over all baselines except LoRA + Camera pose, which we observe to overfit on training images. Figure 6 shows the CLIP vs. DINO scores for all methods and object categories. Ideally, a method should have both a high CLIP score and a DINO score, but often, there is a trade-off between text- and image alignment. Our method has on-par or better text alignment relative to the baselines while having better image alignment. We observe that image-editing baselines often require careful hyperparameter tuning for each image. We select the best-performing hyperparameters and keep them fixed across all experiments. We use the $\sim 50$ validation camera poses not used during training for evaluation and randomly perturb the camera position or focal length. Figure 11 in Appendix B shows sample training and perturbed validation camera poses for the car object.
@@
Camera pose accuracy.   Previously, we have measured our method purely on image customization benchmarks. Next, we evaluate the camera pose accuracy as well. Table 2 shows the camera pose accuracy of the generated images in terms of mean angular error and camera center error. We observe that LoRA + Camera pose baseline overfits on training images and can fail at generating images in the correct pose with new text prompts during inference. We evaluate this on validation camera poses of concepts from the CO3Dv2 dataset with the camera’s principal axis pointing towards the object at the scene’s center. This is because RayDiffusion has been trained on this setup of the CO3Dv2 dataset and fails on unique objects.
@@
Qualitative comparison.   We show the qualitative comparison of our method with the baselines in Figure 4. As we can see, image-editing-based methods often fail at generating photorealistic results. In the case of LoRA + Camera Pose, we observe that it fails to generalize and overfits to the training views ($5^{\text{th}}$ row Figure 4). Finally, the 3D editing-based method Vica-NeRF maintains 3D consistency but generates blurred images for text prompts that change the background scene. Figure 5 shows samples with different text prompts and target camera poses for our method.
@@
Generalization to novel camera poses.   Since our method learns a 3D radiance field, we can also extrapolate to unseen camera poses at inference time as shown in Figure 7. We generate images while varying the camera distance from the object (scale), focal length, or camera position along the horizontal and vertical axis.
@@
Figure 8: Applications. $1^{\text{st}}$ row: Our method can be combined with other image editing methods as well. We use SDEdit with our method to in-paint the rubber duck in different poses while keeping the same background. $2^{\text{nd}}$ row: We can generate interesting panorama shots by controlling the camera pose of the object in each grid independently. $3^{\text{rd}}$ row: We can also compose the radiance field predicted by FeatureNeRF to control the relative pose while generating multiple instances of the object.
Applications.   Our method can be combined with existing image editing methods as well. Figure 8a shows an example where we use SDEdit [52] to in-paint the object in varying poses while keeping the same background.
@@
We can also generate interesting panoramas using MultiDiffusion [3], where the object’s placement in each grid is controlled by our method, as shown in Figure 8b. Moreover, since we learn a 3D consistent FeatureNeRF for the new concept, we can compose multiple instances of the object in feature space [84], with each instance in a different camera pose. Figure 8c shows an example of two teddy bears facing each other sitting on the armchair. Here, we additionally use DenseDiffusion [42] to modulate the attention maps and guide the generation of each object instance to only appear in the corresponding region predicted by FeatureNeRF. At the same time, the attention maps of the empty region predicted by FeatureNeRF are modulated to match the part of the text prompt describing the image’s background.
@@
Table 3: Ablation experiments. Not enriching volumetric features with text cross-attention (Eqn. 6) has an adverse effect on image alignment. Not having mask-based losses (Eqn. 11) leads to overfitting on training images and decreases the text alignment. The worst performing metrics are grayed. Our final method achieves a balance between the input conditions of the target concept, text prompt, and camera pose.
@@ 4.2 Ablation
Background losses.   When removing the silhouette and background loss, as explained in Eqn. 11 from training, we observe a decrease in text alignment and overfitting on training images as shown in Table 3. Figure 12 in Appendix A shows qualitatively that the model generates images with backgrounds more similar to the training views. This is also reflected by the higher similarity between generated images and background regions of the training images ($3^{\text{rd}}$ column Table 3) compared to our final method.
@@
Figure 9: Limitations. Our method can fail when extrapolating camera poses far from the training image camera poses, e.g., changing focal length (top left) or translating camera s.t. the object is not in the center (top right) as the pre-trained model is often biased towards generating the object in the center. Also, it can fail to follow the input text prompt or the exact camera pose when multiple objects are composed in a scene (bottom row).
@@ 5 Discussion and Limitations
We introduce a new task of customizing text-to-image models with camera viewpoint control. Our method jointly learns 3D feature prediction modules and adapts 2D diffusion attention modules to be conditioned on these features. This enables synthesizing the object with new text prompts and precise object pose control. While our method outperforms existing image editing and model customization approaches, it still has several limitations, as we discuss below.
Limitations.   As we show in Figure 9, our method struggles at generalizing to extreme camera poses that were not seen during training and resorts to either changing the object identity or generating the object in a seen pose. We expect this to improve by adding more camera pose variations during training. An additional case when our method struggles to follow the camera pose is when the text prompt adds multiple new objects. We hypothesize that in such challenging scenarios, the model is biased towards generating front views, often seen in training. Also, our proposed pose-conditioning module is trained with a finetuning-based method. Exploring pose-conditioning in a zero-shot customization and editing methods [16, 24] may help reduce the time and computation as well. Here, we focus on enabling camera view control while generating rigid objects. Future work includes extending this conditioning to handle dynamic objects that change the pose in between reference views. One potential way to address this is using a representation based on dynamic and non-rigid NeRF methods [21, 66, 84].
@@
In Appendix A, we show more results and baseline comparisons. In Appendix B, we provide details regarding the evaluation and human preference study. Finally, in Appendix C, we describe all the implementation details of our method and baselines.
Appendix A Results
More baselines.   We add a comparison to SDEdit with the Stable Diffusion-XL model here, which we found to perform worse or on par with the Stable Diffusion-1.5 version of the model. Figure 10 shows the quantitative comparison.

### v2

@@ 3.2 Customization with Object Viewpoint Control
where $M$ is the object mask and $\lambda_{\text{rgb}}$, $\lambda_{\text{bg}}$, and $\lambda_{\text{s}}$ are hyperparameters to control the rendering quality of intermediate images vs. the final denoised image and are fixed across all experiments. We assume access to the object’s mask in the image, which is used to calculate these losses. The losses for FeatureNeRF are averaged across all pose-conditioned transformer layers.
Figure 4: Qualitative comparison. Given a particular target pose, we show the qualitative comparison of our method with (1) Image editing methods SDEdit, InstructPix2Pix, and LEDITS++, which edit a NeRF-rendered image from the input pose, (2) ViCA-NeRF, a 3D editing method that trains a NeRF model for each input prompt, and (3) LoRA + Camera pose, our proposed baseline where we concatenate camera pose information to text embeddings during LoRA fine-tuning. Our method performs on par or better in keeping the target identity and poses while incorporating the new text prompt—e.g., putting a picnic table next to the SUV car ($1^{\text{st}}$ column)—and following multiple text conditions—e.g., turning the chair red and placing it in a white room ($3^{\text{rd}}$ column). V token is used only in ours and the LoRA + Camera pose method. Ground truth rendering from the given pose is shown as an inset in the first three rows. We show more sample comparisons in Figure 17 of Appendix.
Figure 5: Qualitative samples with varying object viewpoint and text prompt. Our method learns the identity of custom objects while allowing the user to control the object viewpoint and generating the object in new contexts using the text prompt, e.g., changing the background scene or object color and shape. In each row, the images were generated with the same seed while changing the object viewpoint in a turntable manner. Note that each image in a row is independently generated. Figure 18 in the Appendix shows more such samples.
@@
Training details.   During training, we sample the $N$ views equidistant from each other and use the first as the target viewpoint and the others as references. We modify $12$ transformer layers with pose conditioning out of $70$ transformer layers in Stable Diffusion-XL. For rendering, we sample $24$ points along the ray. The new concept is described as “V category”, with V as a trainable token embedding [47, 25]. Furthermore, to reduce overfitting [74], we use generated images of the same category, such as random car images with ChatGPT-generated captions [14]. These images are randomly sampled $25\%$ of the time during training. We also drop the text prompt with $10\%$ probability to be able to use classifier-free guidance. We provide more implementation details in Appendix D.
@@ 4 Experiments
Dataset.   We select 14 custom objects from the CO3Dv2 [71] and NAVI [35] datasets. Specifically, we select $4$ categories with $3$ instances each from the CO3Dv2 dataset – car, chair, teddy bear, and motorcycle. From NAVI, we select $2$ unique, toy-like objects. A representative image of each concept is shown in the supplemental material. We use the camera poses provided in the dataset. For each instance, we sample $\sim 100$ images, using half for training and half for evaluation. The camera poses are normalized such that the mean of camera location is the origin, and the first camera is at unit norm [105].
@@
Table 1: Human preference evaluation. Our method is preferred over almost all baselines for text alignment, image alignment to the target concept, and photorealism. We find that LoRA + Camera pose overfits the training images, as shown in Figure 4.
@@
Table 2: Accuracy of object viewpoint condition in generated images by ours and the LoRA + Camera pose baseline method. We observe that the baseline usually overfits the training images and does not respect the target viewpoint condition with new text prompts.
Baselines.   We compare with three types of relevant baselines – (1) First, we compare against 2D image editing using $3$ recent, publicly available methods – LEDITS++ [9], InstructPix2Pix [10], and SDEdit [53] with Stable Diffusion-1.5 (and SDXL in Appendix 4.1). As image editing methods do not inherently support viewpoint manipulation, we first render a NeRF [87] of the input scene with the target viewpoint and then edit the rendered image. (2) Secondly, we use a customization-based method, LoRA+Camera pose, where we modify LoRA [76, 34] by concatenating the camera pose to the text embeddings, following recent work Zero-1-to-3 [50]. (3) Lastly, we test ViCA-NeRF [22], a 3D editing method that trains a NeRF for each new text prompt. In Appendix D, we provide more details on implementation and hyperparameters for each baseline.
Evaluation metrics.   To create an evaluation set, we generate $16$ prompts per object category using ChatGPT [14]. We instruct ChatGPT to propose four types of prompts: scene change, color change, object composition, and shape change. We then manually inspect them to remove implausible or overly complicated text prompts [93]. Table 5 in Appendix C lists all the evaluation prompts. We evaluate (1) the customization quality of the generated image and (2) its adherence to the specific pose.
First, to measure customization quality, we use a pairwise human preference study. A successful customization is comprised of several aspects: alignment to the target concept, alignment to the input text prompt, and photorealism of the generated images. In total, we collect $\sim 1000$ responses per pairwise study against each baseline using Amazon Mechanical Turk. We also evaluate our method and baselines on other standard metrics like CLIP Score [67] and DINOv2 [61] image similarity [74] to measure the text and image alignment.
To measure whether the generated custom object adheres to the specified viewpoint, we use a pre-trained model, RayDiffusion [106], to predict the pose from the generated images and calculate its error relative to the input camera pose. More details about evaluation are provided in Appendix C.
@@ 4.1 Results
Figure 7: Extrapolating object viewpoint from training viewpoints. Our method can generalize to different viewpoints, including those not within the training distribution. Top left: We vary the focal length from ×$\times 0.8$ to ×$\times 1.4$ of the original focal length. Top right: We vary the camera position towards the image plane along the $z$ axis. Bottom row: We vary the camera position along the horizontal and vertical axis.
@@
Generation quality and adherence.   First, we measure the quality of the generation – adherence to the text prompt, the identity preservation to the customized objects, and photorealism – irrespective of the object viewpoint. Recall that for each concept, we curate 16 prompts. For each prompt, we generate $3$ images at each viewpoint, covering $6$ target viewpoints, resulting in $288$ images per concept. Table 1 shows the pairwise human preference for our method vs. baselines. Our method is preferred over all baselines except LoRA + Camera pose, which we observe to overfit on training images, thus producing higher image alignment. Figure 6 shows the CLIP vs. DINO scores for all methods. Ideally, a method should have both a high CLIP score and a DINO score, but often, there is a trade-off between text and image alignment. Our method has on-par or better text alignment relative to the baselines, while having better image alignment. We observe that image-editing baselines often require careful hyperparameter tuning for each image. We select the best-performing hyperparameters and keep them fixed across all experiments. The camera pose corresponding to the target object viewpoint is uniformly sampled from $\sim 50$ validation poses not used during training. We also randomly perturb the camera position or focal length. Figure 13 in Appendix C shows sample training and perturbed validation camera poses for the car object.
@@
Accuracy of object viewpoint.   Previously, we evaluated our method purely on image customization benchmarks. Next, we evaluate the accuracy of the object viewpoint conditioning. Table 2 shows the mean angular error and camera center error between the generated object’s pose, predicted using RayDiffusion [106], and the input pose. We only compare with LoRA + Camera pose, as only this baseline takes the camera pose for the target object viewpoint as input. We observe that it often overfits training images and fails to generate the object in the correct viewpoint with new text prompts. We evaluate this metric only on the objects from the CO3Dv2 dataset with validation camera poses, as RayDiffusion has been trained on CO3Dv2 and struggles with other unique objects.
@@
Qualitative comparison.   We show the qualitative comparison of our method with the baselines in Figure 4. We observe that image editing methods can fail to generate photorealistic results. In the case of LoRA + Camera pose, it fails to generalize and overfits to the training views ($5^{\text{th}}$ row). Finally, the 3D editing-based method ViCA-NeRF maintains 3D consistency but generates blurred images, especially for text prompts that change the background. Figure 5 shows more samples with different text prompts and object viewpoints for our method.
Additional comparison to customization + 3D-aware image editing.   We further compare against a two-stage approach that first generates an image of the custom object using LoRA+DreamBooth [76, 74] and then edits the object to a target viewpoint using two recent 3D-aware image editing methods, Image Sculpting [101] and Object3DIT [55]. For each prompt, we generate $3$ images, then edit and rotate the object to 6 different viewpoints. This results in 288 images per concept, similar to our evaluation setting. We compare against this on only the three car objects since Image-Sculpting uses Adobe Photoshop’s generative fill [1] as one of the intermediate steps, which requires manually inpainting each image. The CLIP scores for Image Sculpting and Object3DIT are $0.26$ and $0.27$, respectively, compared to our score of $0.25$. However, their DINO scores at $0.24$ and $0.40$ are substantially lower than our $0.48$. As shown in Figure 8, both methods lead to lower-fidelity results. Object3DIT struggles in many scenarios due to its training on a synthetic dataset, and Image Sculpting’s performance is highly dependent on single image-to-3D methods like Zero-1-to-3 [50] used in its pipeline.
@@
Generalization to novel viewpoints.   Since our method learns a 3D radiance field, we can also extrapolate to unseen object viewpoints at inference time as shown in Figure 7. We generate images while varying the camera distance from the object (scale), focal length, or camera position along the horizontal and vertical axis.
Figure 8: Comparison to 3D-aware editing methods. We first generate the image of the custom object using LoRA+DreamBooth (shown as an inset) and then use the 3D-aware editing method to edit and rotate the object to a target viewpoint. We show qualitative samples generated by our method ($3^{\text{rd}}$ column) with approximately the same target viewpoint as input. Object3DIT and Image-Sculpting lead to lower fidelity edits than images generated by our method ($3^{\text{rd}}$ column) with the target viewpoint directly as the input condition.
@@
Figure 9: Applications. $1^{\text{st}}$ row: Our method can be combined with other image editing methods as well. We use SDEdit with our method to in-paint the car and rubber duck from different viewpoints while keeping the same background. $2^{\text{nd}}$ row: We can generate interesting panorama shots by controlling the object viewpoint independently in each grid. $3^{\text{rd}}$ row: We can also compose the radiance field predicted by FeatureNeRF to control the relative pose while generating multiple instances of the object.
@@
Applications.   Our method can be combined with existing image editing methods as well. Figure 9a shows an example where we use SDEdit [53] to in-paint the object with different viewpoints while keeping the same background. We can also generate interesting panoramas using MultiDiffusion [4], where the object viewpoint in each grid is controlled by our method, as shown in Figure 9b. Moreover, since we learn a 3D consistent FeatureNeRF for the new object, we can compose multiple instances of the object [84], with each instance in a different viewpoint. Figure 9c shows an example of two teddy bears facing each other and sitting on armchairs. Here, we additionally use DenseDiffusion [44] to modulate the attention maps and guide the generation of each object instance to only appear in the corresponding region predicted by FeatureNeRF. At the same time, the attention maps of the empty region predicted by FeatureNeRF are modulated to match the part of the text prompt describing the image’s background.
@@
Table 3: Ablation experiments. Not enriching volumetric features with text cross-attention (Eqn. 6) has an adverse effect on image alignment. Not having mask-based losses (Eqn. 10) leads to overfitting on training images and decreases the text alignment. The worst performing metrics are grayed. Our final method achieves a balance between the input conditions of the target concept, text prompt, and camera pose.
@@ 4.2 Ablation
Background losses.   When removing the silhouette and background loss, as explained in Eqn. 10 from training, we observe a decrease in text alignment and overfitting on training images as shown in Table 3. Figure 14 in Appendix 4.1 shows qualitatively that the model generates images with backgrounds more similar to the training views. This is also reflected by the higher similarity between generated images and background regions of the training images ($3^{\text{rd}}$ column Table 3) compared to our final method.
@@
Figure 10: Limitations. Our method can occasionally fail when the target object viewpoint deviates far from the training images, e.g., reducing the focal length too much (top left) or rendering the object off-center (top right), as the pre-trained model is often biased towards generating the object in the center. Also, it can fail to follow the input text prompt or the exact object viewpoint when multiple objects are composed in a scene (bottom row).
We show more results and ablation experiments in the Appendix, including performance with predicted camera viewpoints.
@@ 5 Discussion and Limitations
We introduce a new task of customizing text-to-image models with object viewpoint control. Our method learns view-dependent object features in the intermediate feature space of the diffusion model and conditions the generation on them. This enables synthesizing the object with varying object viewpoints while controlling other aspects through text prompts.
Limitations.
Though our method outperforms existing image editing and model customization approaches, it still has several limitations. As we show in Figure 10, our method occasionally struggles at generalizing to extreme viewpoints that were not seen during training and resorts to either changing the object identity or generating the object in a seen viewpoint. We expect this to improve by adding more viewpoint variations during training. Our method also sometimes struggles to follow the input viewpoint condition when the text prompt adds multiple objects to the scene. We hypothesize that in such challenging scenarios, the model is biased towards generating object-centric front views, as seen in its original training data. Also, we fine-tune the model for each custom object, which takes computation time ($\sim$ 40 minutes). Exploring pose-conditioning in a zero-shot, feed-forward manner [18, 26] may help reduce the time and computation. Finally, we focus on enabling viewpoint control for rigid objects. Future work includes extending this conditioning to handle dynamic objects that change the pose in between reference views. One potential way to address this is using a representation based on dynamic and non-rigid NeRF methods [23, 66, 84].
@@
In Appendix A and Section B, we show more ablation and results. In Appendix C, we provide details regarding the evaluation and human preference study. Finally, in Appendix D, we describe all the implementation details of our method and baselines.
Appendix A Ablation
Using predicted masks for background losses.   In our main paper, we calculated the background losses (Eqn. 10) using ground truth masks from the dataset during training. We show in Table 4 that replacing these with predicted masks results in similar performance. We use Detic [110] with Segment Anything [46] to predict the object mask given the object category, such as car and teddy bear.
Using predicted camera viewpoints.   Here, we use COLMAP to predict the camera viewpoints using only the $50$ training images. Though it requires longer training, $2000$ steps compared to $1600$ for our method, the final performance is comparable to our original method. The CLIP and DINO scores remain similar while the mean angular error increases slightly to $17.12$ compared to $16.14$ of our final method on $10$ objects (COLMAP fails to run on $2$ car objects). Qualitative samples and their comparison to our final method are shown in Figure 11.
Appendix B More Results